OpenStudy (anonymous):

32,use the Newton-Raphson method to estimate a root of the equation f (x) = 0 starting at the indicated value of x:(a) Express xn+1 in terms of xn. (b) Give x4 rounded off to five decimal places and evaluate f at that approximation f (x) = x^2 − 24; x1 = 5.

OpenStudy (anonymous):

im real confused on how to go about solving this problem

OpenStudy (turingtest):

\[x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}\]

OpenStudy (turingtest):

for n=1 what does that formula give?

OpenStudy (anonymous):

xn+1=1/2xn+12(1/xn)

OpenStudy (turingtest):

I have no idea what you just did start at \(x_1=5\), we know \(f(x)=x^2-24\) so what is? \[x_2=x_1-\frac{f(x_1)}{f'(x_1)}\]

OpenStudy (anonymous):

5-1/10?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

that is x2 now repeat that 2 more times to get x4, which is what the problem requests